An chart with three-sigma limits has parameters a
Answer #1

Ok now I am going to solve this BIG question just for sake of brotherhood of industrial Engineers \m/


Now amount of output falling outside of Control Limits.

Above z=(104-98)/8/root(5) = 1.677

amount of output falling outside of Upper Control Limits =1-P(z=1.67) = 5% from normal tables


below Z= (96-98)/8/root(5) = -0.56 = P(z=-0.56) = 28.7%

amount of output falling outside of Lower Control Limits =28.7%


totaloutput falling outside of Control Limits =33.7%

so Probability that CL would exibit lack of control for at least third poit plotted = 0.663*0.663*0.337 = 0.148 = 14.8%



Please Rate 5 star for help and Explaination


.